OpenStudy (anonymous):

(x-2)^3/2=125 solution set?

OpenStudy (anonymous):

Is the problem: \[(x - 2)^{\frac32} = 125\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright. \[(x - 2)^{\frac32} = 125\] \[((x - 2)^{\frac32})^{\frac23} = (125)^{\frac23}\] \[x - 2 = 125^{\frac23}\] \[x - 2 = (\sqrt[3]{125})^{2}\] Can you solve it now?

OpenStudy (anonymous):

hopefully

OpenStudy (anonymous):

Let me know what you get.

OpenStudy (anonymous):

27?

OpenStudy (anonymous):

There you go :)

OpenStudy (anonymous):

so would it be 27 and -27

OpenStudy (anonymous):

for the solution set

OpenStudy (anonymous):

Is this multiple choice?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Alright. I forgot to put this in, so just watch closely. \[x - 2 = ±(\sqrt[3]{125})^{2}\] \[x - 2 = ±25\]

OpenStudy (anonymous):

ohh got it

OpenStudy (anonymous):

Alright. One of the answers is 27, but what do you get for the second one?

OpenStudy (anonymous):

-23

OpenStudy (anonymous):

There you go :). That should work.

OpenStudy (anonymous):

what about logsmall3 3 ^8